OpenStudy (anonymous):

Find f(6) if f(x) = x2 ÷ 3 + x.

OpenStudy (anonymous):

f(x) = x2 ÷ 3 + x. f(6) = 62 ÷ 3 + 6. solve it now

OpenStudy (anonymous):

The answer has to be one of these following numbers.. 4, 10, or 18.